# Diagonal comparison of Ample $\mathrm{C}^{\ast}$-diagonals

### v1

@@ 1. Diagonal comparison
Notice that $\precsim_{\mathrm{diag}}$ is stronger than the usual Cuntz subequivalence, since for $a\precsim_{\mathrm{diag}}b$ to hold we are asking that $a$ is Cuntz below $b$ via elements that factorise as a product of three elements as described above. Also notice that the relation $\precsim_{\mathrm{diag}}$ is not designed to be transitive.
@@
Let $p,q\in D$ be projections with $\tau(p)<\tau(q)$ for all $\tau\in\mathrm{T}(A)$. By diagonal comparison we have $p\precsim_{\mathrm{diag}}q$, so by Lemma 1.12 there are $u\in\mathrm{her}(p)$, $v\in\mathrm{her}(q)$ and $w\in\mathcal{N}_{A}(D)$ such that
$\|p-uwvqv^{*}w^{*}u^{*}\|<1.$
In particular, $uwvqv^{*}w^{*}u^{*}\in\mathrm{her}(p)$ is a positive element that is invertible in $\mathrm{her}(p)$, and, since $vqv^{*}\leq\|v\|^{2}q$, so are $uwqw^{*}u^{*}$ and $uu^{*}$. Since $\mathrm{her}(p)$ is stably finite as a hereditary subalgebra of $A$, right invertibility of $u$ implies that $u$ itself is invertible in $\mathrm{her}(p)$, and thus so is $pwqw^{*}p$. Now $D$ is abelian and by invertibility of $pwqw^{*}$ in $\mathrm{her}(p)$ it follows that $wqw^{*}\in D$ does not vanish on the support of $p$ which is a clopen set, whence there is $h\in D$ such that $hwqw^{*}h=p$. Setting $\tilde{w}:=hw\in\mathcal{N}_{A}(D)$ we have that $p=\tilde{w}q\tilde{w}^{*}$, proving that $p\prec_{(D\subset A)}q$ as we wanted. ∎
@@ 3. Main result and consequences
Let $n\in\mathbb{N}$ and let $\{e_{i,j}\}_{i,j=1}^{n}$ be the canonical system of matrix units for $M_{n}$. For any $d=\sum_{i=1}^{n}d_{i}\otimes e_{i,i}\in D\otimes D_{n}$ with $d_{i}\in D_{+}$, $i=1,\dots,n$, a direct argument similar to the one above shows that
@@
If $A$ is finite-dimensional, then it is a matrix algebra due to simplicity, and diagonal comparison can be verified directly: if $a,b$ are positive matrices with the rank of $a$ being less than the rank of $b$ and $p_{a},p_{b}$ are the (diagonal) support projections of the conditional expectations $E(a)$ and $E(b)$ respectively, then there is a permutation matrix $w$ (thus a normalizer of diagonal matrices) conjugating $p_{b}$ to a diagonal projection greater than $p_{a}$. The spectrum of $b$ is finite, so $b^{1/2}$ is locally invertible, and thus $b^{1/2}$ is also invertible in $\mathrm{her}(p_{b})$. If $v\in\mathrm{her}(p_{b})_{+}$ is the (local) inverse of $b^{1/2}$, we have that $a=a^{1/2}p_{a}wvbv^{*}w^{*}p_{a}a^{1/2}$, i.e. $a\precsim_{\mathrm{diag}}b$. We can thus restrict to the infinite-dimensional case, which is the nontrivial one.
@@
(i)$\implies$(ii). If $(C(X)\subset C(X)\rtimes_{\mathrm{r}}G)$ has finite diagonal dimension, then by [19, Theorem 5.4] and [14, Theorem 7.2] $G\curvearrowright X$ has dynamical comparison and, since $X$ is zero-dimensional, by [15, Theorem 6.1] $G\curvearrowright X$ is almost finite, so (ii) follows by [21, Corollary 3.7]. (ii)$\iff$(iii) is [21, Corollary 3.7]. (iii)$\iff$(iv) follows from Theorem 3.6 which is applicable since the conditional expectation $C(X)\rtimes_{\mathrm{r}}G\to C(X)$ is hereditary (see Remark 1.10), together with the fact that dynamical comparison of $G\curvearrowright X$ and zero-dimensionality of $X$ imply strict comparison of $C(X)\rtimes_{\mathrm{r}}G$ by [14, Theorem 12.4].
@@ 4. Hereditary conditional expectations
Let $r_{\max}\in\mathbb{N}$ be the maximum matrix size among the matrix summands of $F$. By Urysohn’s lemma there is $f\in D_{+}^{1}$ such that $\mathrm{supp}(f)\subset\mathrm{supp}(\Phi(a))$ and $f=1$ on the subset of the spectrum of $D$ where $\Phi(a)\geq\frac{\eta}{2(d+1)r_{\max}}$, in particular $f\in\mathrm{her}(\Phi(a))$. Set $f^{\prime}\coloneqq 1-f\in D_{+}^{1}$. Note that $a=faf+f^{\prime}af+faf^{\prime}+f^{\prime}af^{\prime}$ and therefore
@@
whence, by our choice of $\eta$, it suffices to show that $\|f^{\prime}\varphi\psi(a)f^{\prime}\|<\eta$. Setting $\psi^{(j)}\coloneqq\pi_{F^{(j)}}\circ\psi$ where $\pi_{F^{(j)}}$ is the canonical surjection $F\to F^{(j)}$, since $\varphi\psi=\sum_{j=0}^{d}\varphi^{(j)}\psi^{(j)}$, it further suffices to show that $\|f^{\prime}\varphi^{(j)}\psi^{(j)}(a)f^{\prime}\|<\frac{\eta}{d+1}$ for all $j=0,\dots,d$, since then it follows by the triangle inequality that
@@
Fix $j_{0}\in\{0,\dots,d\}$ and set $B\coloneqq\mathrm{C}^{*}(\varphi^{(j_{0})}(F^{(j_{0})}),D)\subset A$. For a matrix summand $E$ appearing in $F^{(j_{0})}$ let $J_{E}:=\langle\varphi_{E}^{(j_{0})}(E)\rangle\trianglelefteq B$ be the ideal generated by $\varphi_{E}^{(j_{0})}(E)$ in $B$. If $E,K$ are distinct matrix summands appearing in $F^{(j_{0})}$, then straightforward calculations show that $J_{E}J_{K}=\{0\}$.
Since $f^{\prime}\varphi^{(j_{0})}\psi^{(j_{0})}(a)f^{\prime}=\sum_{E}f^{\prime}\varphi_{E}^{(j_{0})}\psi^{(j_{0})}_{E}(a)f^{\prime}$, and this is a sum of pairwise orthogonal contractions, we have that
@@
Let $E=M_{r}$ be a matrix summand appearing in $F^{(j_{0})}$ and set $h_{E}\coloneqq f^{\prime}\varphi_{E}^{(j_{0})}\psi_{E}^{(j_{0})}(a)f^{\prime}\in J_{E}$. To simplify notation, set $\varphi_{E}\coloneqq\varphi_{E}^{(j_{0})}$. As $E$ is simple, the structure theorem of order zero maps [36, Theorem 2.3] yields that $\varphi_{E}$ is either zero or injective. In the first case $h_{E}=0$ and there is nothing to do, so assume that $\varphi_{E}$ is injective. Letting $\{e_{i,j}\}_{i,j=1}^{r}$ be the canonical matrix units of $E$, it is not hard to see that $J_{E}\cap\mathrm{her}(\varphi_{E}(e_{i,i}))\subset D$ for all $i=1,\dots,r$.
Let $\pi\colon J_{E}\to\mathbb{B}(\mathcal{H})$ be an irreducible representation of $J_{E}$ such that $\|h_{E}\|=\|\pi(h_{E})\|$. For $i=1,\dots,r$ consider the limits in the strong operator topology
$p_{i}:=\mathrm{s.o.}\lim_{n\to\infty}\pi(\varphi_{E}(e_{i,i})^{1/n})\in\mathbb{B}(\mathcal{H}).$
Then $\{p_{i}\}_{i=1}^{r}$ are pairwise orthogonal projections and $\mathrm{id}_{\mathcal{H}}=\sum_{i=1}^{r}p_{i}$, since $\{\varphi_{E}(1_{E})^{1/n}\}_{n=1}^{\infty}$ is an approximate unit for $J_{E}$. Moreover, it is easy to see that for any $x\in J_{E}\cap\mathrm{her}(\varphi_{E}(e_{i,i}))$ one has $\pi(x)=p_{i}\pi(x)=\pi(x)p_{i}$, and therefore since each $J_{E}\cap\mathrm{her}(\varphi_{E}(e_{i,i}))$ is abelian and non-zero, each projection $p_{i}$ has rank one and thus $\mathcal{H}$ is $r$-dimensional, whence $\mathbb{B}(\mathcal{H})$ can be identified with $M_{r}$ in such a way so that $\pi(\varphi_{E}(e_{i,i}))$ is a scalar multiple of $e_{i,i}$ for each $i=1,\dots,r$. Therefore, the composition $\pi\circ\varphi_{E}\colon M_{r}\to M_{r}$ is a bijective c.p.c. order zero map and maps $e_{i,i}$ to a scalar multiple of itself. It follows by the structure theorem that there is $\lambda>0$ such that
@@
Let $\sigma\colon J_{E}\cap\mathrm{her}(\varphi_{E}(e_{1,1}))\to\mathbb{C}$ be the character satisfying $\pi(x)=\sigma(x)\cdot e_{1,1}$ for all $x\in J_{E}\cap\mathrm{her}(\varphi_{E}(e_{1,1}))$. Since $J_{E}\cap\mathrm{her}(\varphi_{E}(e_{1,1}))\subset D$, we can extend $\sigma$ to a character on $D$ which we denote again by $\sigma$, and which we can in turn extend to a state on $A$ as $\tilde{\sigma}:=\sigma\circ\Phi$. Define a map $\varrho\colon A\to M_{r}$ as
@@
We now claim that, for $g\in D_{+}$ and $e_{i,j}\in M_{r}$ we have $\varrho(g\varphi_{E}(e_{i,j})g)=\pi(g\varphi_{E}(e_{i,j})g)$. Indeed, using the fact that $\varphi_{E}(D_{r})$ is a subset of $D$ which is abelian, by order zero functional calculus [36, Corollary 3.2] we see that $g\varphi_{E}(e_{i,j})g\in\overline{\varphi_{E}(e_{i,i})A\varphi_{E}(e_{j,j})}$ and therefore
@@
which in conjuction with (4.5) shows that $\varrho(g\varphi_{E}(e_{i,j})g)=\pi(g\varphi_{E}(e_{i,j})g)$ as we wanted. It follows now that $\varrho(h_{E})=\pi(h_{E})$. Since $\pi$ was picked so that $\|h_{E}\|=\|\pi(h_{E})\|$, we have that $\|h_{E}\|=\|\varrho(h_{E})\|$. Let $h^{\prime}_{E}:=f^{\prime}a^{1/2}\varphi_{E}^{(j_{0})}\psi_{E}^{(j_{0})}(1_{A})a^{1/2}f^{\prime}$. Since $\|h_{E}-h^{\prime}_{E}\|<\frac{\eta}{2(d+1)}$ by (vi) of Proposition 4.1 and since $\varrho$ is contractive, we have that
@@
We also have that

### v2

@@ 1. Diagonal comparison
Notice that $\precsim_{\mathrm{diag}}$ is stronger than the usual Cuntz subequivalence, since for $a\precsim_{\mathrm{diag}}b$ to hold we are asking that $a$ is Cuntz below $b$ via elements that factorise as a product of three elements as described above.
@@
Let $p,q\in D$ be projections with $\tau(p)<\tau(q)$ for all $\tau\in\mathrm{T}(A)$. If $p=0$ there is nothing to show, so we assume that $p\neq 0$. By diagonal comparison we have $p\precsim_{\mathrm{diag}}q$, so by Lemma 1.12 there are $u\in\mathrm{her}(p)$, $v\in\mathrm{her}(q)$ and $w\in\mathcal{N}_{A}(D)$ such that
$\|p-uwvqv^{*}w^{*}u^{*}\|<1,$
In particular, $u,w,v$ are non-zero, and $uwvqv^{*}w^{*}u^{*}\in\mathrm{her}(p)$ is a positive element that is invertible in $\mathrm{her}(p)$. Moreover, since $v\in\mathrm{her}(q)$, we have $vqv^{*}\leq\|v\|^{2}q$, and so $uwqw^{*}u^{*}$ is invertible in $\mathrm{her}(p)$, and since $0\leq uwqw^{*}u^{*}\leq\|w\|^{2}uu^{*}$, $uu^{*}$ is also invertible in $\mathrm{her}(p)$. Since $\mathrm{her}(p)$ is stably finite as a hereditary subalgebra of $A$, right invertibility of $u$ implies that $u$ itself is invertible in $\mathrm{her}(p)$, and thus so is $pwqw^{*}p$. Now $D$ is abelian and by invertibility of $pwqw^{*}p$ in $\mathrm{her}(p)$ it follows that $wqw^{*}\in D$ does not vanish on the support of $p$ which is a clopen set, whence there is $h\in D$ such that $hwqw^{*}h=p$. Setting $\tilde{w}:=hw\in\mathcal{N}_{A}(D)$ we have that $p=\tilde{w}q\tilde{w}^{*}$, proving that $p\prec_{(D\subset A)}q$ as we wanted. ∎
@@
Remark 1.14.
We note that the relation $\precsim_{\mathrm{diag}}$ is not designed to be a priori transitive. However, in the presence of diagonal comparison, the combination of $\precsim_{\mathrm{diag}}$ with strict subequality on all dimension functions, indeed yields a transitive relation. It remains an interesting technical problem whether the relation $\precsim_{\mathrm{diag}}$ alone can fail to be transitive.
@@ 3. Main result and consequences
Let $n\in\mathbb{N}$ and let $\{e_{i,j}\}_{i,j=1}^{n}$ be the canonical system of matrix units for $M_{n}$. For any $d=\sum_{i=1}^{n}d_{i}\otimes e_{i,i}\in D\otimes D_{n}$ with $d_{i}\in D_{+}$, $i=1,\dots,n$, one directly sees that
@@
If $A$ is finite-dimensional, then it is a matrix algebra due to simplicity, and diagonal comparison can be verified directly: if $a,b$ are positive matrices with the rank of $a$ being less than the rank of $b$ and $p_{a},p_{b}$ are the (diagonal) support projections of the conditional expectations $E(a)$ and $E(b)$ respectively, then there is a permutation matrix $w$ (thus a normaliser of diagonal matrices) conjugating $p_{b}$ to a diagonal projection greater than $p_{a}$. The spectrum of $b$ is finite, so $b^{1/2}$ is locally invertible, and thus $b^{1/2}$ is also invertible in $\mathrm{her}(p_{b})$. If $v\in\mathrm{her}(p_{b})_{+}$ is the (local) inverse of $b^{1/2}$, we have that $a=a^{1/2}p_{a}wvbv^{*}w^{*}p_{a}a^{1/2}$, i.e. $a\precsim_{\mathrm{diag}}b$. We can thus restrict to the infinite-dimensional case, which is the nontrivial one.
@@
(ii)$\iff$(iii) is [21, Corollary 3.7]. The equivalence (iii)$\iff$(iv) follows from Theorem 3.6 which is applicable since the conditional expectation $C(X)\rtimes_{\mathrm{r}}G\to C(X)$ is hereditary (see Remark 1.10), together with the fact that dynamical comparison of $G\curvearrowright X$ and zero-dimensionality of $X$ imply strict comparison of $C(X)\rtimes_{\mathrm{r}}G$ by [14, Theorem 12.4]. Finally, if $(C(X)\subset C(X)\rtimes_{\mathrm{r}}G)$ has finite diagonal dimension, then by [19, Theorem 5.4] and [14, Theorem 7.2] $G\curvearrowright X$ has dynamical comparison, so (i)$\implies$(iii).
@@ 4. Hereditary conditional expectations
Remark 4.2.
For the proof of Proposition 4.3 below we will need that order zero maps and order zero functional calculus are compatible with preservation of normalisers: if $(D_{A}\subset A)$ and $(D_{B}\subset B)$ are unital sub-$\mathrm{C}^{\ast}$-algebras with $D_{A},D_{B}$ abelian and $\varphi\colon A\to B$ is a c.p.c. order zero map that preserves normalisers, then we automatically have $\varphi(D_{A})\subset D_{B}$ (cf. [19, Proposition 1.9 (i)] or [21, Lemma 1.6]). Moreover, for a positive $f\in C_{0}(0,1]$, the c.p.c. order zero map $f(\varphi)\colon A\to B$ defined by order zero functional calculus (see [36, Corollary 3.2]) preserves normalisers: indeed, let $\pi_{\varphi}$ denote the -homomorphism associated to $\varphi$ by the structure theorem of order zero maps [36, Theorem 2.3], so $\varphi(.)=\varphi(1_{A})\pi_{\varphi}(.)$. For $v\in\mathcal{N}_{A}(D_{A})$, we have $f(\varphi)(v)=f(\varphi(1_{A}))\cdot\pi_{\varphi}(v)$. Uniformly approximating $f$ by polynomials with zero constant term, we can find a sequence $(d_{n})_{n=1}^{\infty}\subset D_{B}$ such that $f(\varphi(1_{A}))=\lim_{n\to\infty}d_{n}\cdot\varphi(1_{A})$, whence $f(\varphi)(v)=\lim_{n\to\infty}d_{n}\varphi(1_{A})\pi_{\varphi}(v)=\lim_{n\to\infty}d_{n}\varphi(v)\in\mathcal{N}_{B}(D_{B})$.
@@
Let $r_{\max}\in\mathbb{N}$ be the maximum matrix size among the matrix summands of $F$. By Urysohn’s lemma there is $f\in D_{+}^{1}$ such that $\mathrm{supp}(f)\subset\mathrm{supp}(\Phi(a))$ and $f=1$ on the subset of the spectrum of $D$ where $\Phi(a)\geq\frac{\eta}{2(d+1)r_{\max}}$, and in particular $f\in\mathrm{her}(\Phi(a))$. Set $f^{\prime}\coloneqq 1-f\in D_{+}^{1}$. Note that $a=faf+f^{\prime}af+faf^{\prime}+f^{\prime}af^{\prime}$ and therefore
@@
whence, by our choice of $\eta$, it suffices to show that $\|f^{\prime}\varphi\psi(a)f^{\prime}\|<\eta$, since then we also have $\|f^{\prime}af^{\prime}\|<2\eta$, and thus $\|a-faf\|<2\eta+2\sqrt{2\eta}<\varepsilon$. Setting $\psi^{(j)}\coloneqq\pi_{F^{(j)}}\circ\psi$ where $\pi_{F^{(j)}}$ is the canonical surjection $F\to F^{(j)}$, since $\varphi\psi=\sum_{j=0}^{d}\varphi^{(j)}\psi^{(j)}$, it further suffices to show that $\|f^{\prime}\varphi^{(j)}\psi^{(j)}(a)f^{\prime}\|<\frac{\eta}{d+1}$ for all $j=0,\dots,d$, since then it follows by the triangle inequality that
@@
Fix $j_{0}\in\{0,\dots,d\}$ and set $B\coloneqq\mathrm{C}^{*}(\varphi^{(j_{0})}(F^{(j_{0})}),D)\subset A$. For a matrix summand $E$ appearing in $F^{(j_{0})}$ let $J_{E}:=\langle\varphi_{E}^{(j_{0})}(E)\rangle\trianglelefteq B$ be the ideal generated by $\varphi_{E}^{(j_{0})}(E)$ in $B$. (These ideals as well as $B$ itself are very well-behaved since $\varphi^{(j_{0})}$ is compatible with $D$; in fact, $B$ and the ideals $J_{E}$ are subhomogeneous, as we will see below.)
If $E,K$ are distinct matrix summands appearing in $F^{(j_{0})}$, then some calculations show that $J_{E}J_{K}=\{0\}$: to see this, by definition of $B$ and since $1_{A}\in D$, it suffices to show that for any $n\geq 0$, any matrix summands $E_{1},\dots,E_{n}$ of $F^{(j_{0})}$, any matrix units $e_{1}\in E_{1},\dots,e_{n}\in E_{n}$, any matrix units $e\in E,e^{\prime}\in K$, and any $f_{1},\dots,f_{n+1}\in D$, we have that
(4.1) $z\coloneqq\varphi^{(j_{0})}(e)\cdot\bigg{(}f_{1}\varphi^{(j_{0})}(e_{1})f_{2}\varphi^{(j_{0})}(e_{2})\cdots f_{n}\varphi^{(j_{0})}(e_{n})f_{n+1}\bigg{)}\cdot\varphi^{(j_{0})}(e^{\prime})=0,$
where, if $n=0$, the middle product is to be interpreted as only $f_{n+1}$. Note that since matrix units of any matrix summand in $F^{(j_{0})}$ are normalisers of $D_{F^{(j_{0})}}$, by (v) of our application of Proposition 4.1 we have that each of the $\varphi^{(j_{0})}(e_{i})$ and $\varphi^{(j_{0})}(e),\varphi^{(j_{0})}(e^{\prime})$ are in $\mathcal{N}_{A}(D)$. Repeatedly using this observation together with the fact that $D$ is commutative, one sees that for $zz^{*}$ to be non-zero it is necessary that $K=E_{n}=\dots=E_{1}=E$, which is impossible.
Since $f^{\prime}\varphi^{(j_{0})}\psi^{(j_{0})}(a)f^{\prime}=\sum_{E}f^{\prime}\varphi_{E}^{(j_{0})}\psi^{(j_{0})}_{E}(a)f^{\prime}$ which is a sum of pairwise orthogonal contractions, we have that
@@
Let $E=M_{r}$ be a matrix summand appearing in $F^{(j_{0})}$ and set $h_{E}\coloneqq f^{\prime}\varphi_{E}^{(j_{0})}\psi_{E}^{(j_{0})}(a)f^{\prime}\in J_{E}$. To simplify notation, set $\varphi_{E}\coloneqq\varphi_{E}^{(j_{0})}$. As $E$ is simple, the structure theorem of order zero maps yields that $\varphi_{E}$ is either zero or injective. In the first case $h_{E}=0$ and there is nothing to do, so assume that $\varphi_{E}$ is injective.
If $\{e_{i,j}\}_{i,j=1}^{r}$ are the matrix units of $E$, we see that $\overline{\varphi_{E}(e_{i,i})B\varphi_{E}(e_{i,i})}\subset D$ for all $i=1,\dots,r$. Indeed, similarly to (4.1), it suffices to show that for any $n\geq 0$, any matrix units $e_{i_{1},j_{1}},\dots,e_{i_{n},j_{n}}$ of matrix summands of $F^{(j_{0})}$, and any $f_{1},\dots,f_{n+1}\in D$, we have that
(4.3) $\varphi_{E}(e_{i,i})\cdot\bigg{(}f_{1}\varphi^{(j_{0})}(e_{i_{1},j_{1}})f_{2}\varphi^{(j_{0})}(e_{i_{2},j_{2}})\cdots f_{n}\varphi^{(j_{0})}(e_{i_{n},j_{n}})f_{n+1}\bigg{)}\cdot\varphi_{E}(e_{i,i})\in D,$
where if $n=0$, the middle product is to be interpreted as only $f_{n+1}$ (and in which case (4.3) is obvious). Let $w$ denote the product in (4.3). If $w$ is non-zero, then all of $\varphi_{E}(e_{i,i})f_{1}\varphi^{(j_{0})}(e_{i_{1},j_{1}})$, $\varphi^{(j_{0})}(e_{i_{1},j_{1}})f_{2}\varphi^{(j_{0})}(e_{i_{2},j_{2}})$, $\dots$, $\varphi^{(j_{0})}(e_{i_{n-1},j_{n-1}})f_{n}\varphi^{(j_{0})}(e_{i_{n},j_{n}})$, $\varphi^{(j_{0})}(e_{i_{n},j_{n}})f_{n+1}\varphi_{E}(e_{i,i})$ are non-zero. By repeated applications of the $\mathrm{C}^{\ast}$-identity, the fact that all of the $\varphi^{(j_{0})}(e_{i_{k},j_{k}})$ are normalisers of $D$ and since $\varphi^{(j_{0})}$ is order zero, it follows that each $e_{i_{k},j_{k}}$ is a matrix unit of $E$, and moreover $i_{1}=i$, $i_{k+1}=j_{k}$ for all $k=1,\dots,n-1$, and $j_{n}=i$. Now using order zero functional calculus, set $\beta\coloneqq(\varphi^{(j_{0})})^{\frac{1}{2}}$, and observe that $\varphi^{(j_{0})}(xy)=\beta(x)\beta(y)$ for all $x,y\in F^{(j_{0})}$. Putting these together, and since $\beta$ preserves normalisers by Remark 4.2, we see that
$\varphi^{(j_{0})}(e_{i_{1},j_{1}})f_{2}\varphi^{(j_{0})}(e_{i_{2},j_{2}})\cdots\varphi^{(j_{0})}(e_{i_{n-1},j_{n-1}})f_{n}\varphi^{(j_{0})}(e_{i_{n},j_{n}})$ $=$
$\varphi^{(j_{0})}(e_{i,j_{1}})f_{2}\varphi^{(j_{0})}(e_{j_{1},j_{2}})\cdots\varphi^{(j_{0})}(e_{j_{n-2},j_{n-1}})f_{n}\varphi^{(j_{0})}(e_{j_{n-1},i})$ $=$
$\underbrace{\beta(e_{i,i})}_{\in D}\cdot\underbrace{\beta(e_{i,j_{1}})f_{2}\beta(e_{j_{1},i})}_{\in D}\cdot\underbrace{\beta(e_{i,j_{2}})f_{3}\beta(e_{j_{2},i})}_{\in D}\cdots$
$\underbrace{\cdots\beta(e_{j_{n-2},i})}_{\in D}\cdot\underbrace{\beta(e_{i,j_{n-1}})f_{n}\beta(e_{j_{n-1},i})}_{\in D}\cdot\underbrace{\beta(e_{i,i})}_{\in D},$
and therefore (4.3) shows that $w\in D$.
Let $\pi\colon J_{E}\to\mathbb{B}(\mathcal{H})$ be an irreducible representation of $J_{E}$ such that $\|h_{E}\|=\|\pi(h_{E})\|$. Note that $\varphi_{E}(1_{E})$ commutes with $B$ (due to the structure theorem of order zero maps and the fact that $\varphi_{E}(1_{E})\in D$), and in particular $\pi(\varphi_{E}(1_{E}))\in\pi(J_{E})^{\prime}=\mathbb{C}\mathrm{id}_{\mathcal{H}}$, so there is $\lambda>0$ such that $\pi(\varphi_{E}(1_{E}))=\lambda\cdot\mathrm{id}_{\mathcal{H}}$. Now $\theta\coloneqq\frac{1}{\lambda}\cdot\pi\circ\varphi_{E}$ is a unital order zero c.p. map, whence it is a unital -homomorphism $M_{r}\to\mathbb{B}(\mathcal{H})$, due to the structure theorem of order zero maps. The projections $p_{i}\coloneqq\theta(e_{i,i})$ are pairwise orthogonal and satisfy $\sum_{i=1}^{r}p_{i}=\mathrm{id}_{\mathcal{H}}$, and for $x\in\overline{\varphi_{E}(e_{i,i})J_{E}\varphi_{E}(e_{i,i})}$ one has $\pi(x)=p_{i}\pi(x)=\pi(x)p_{i}$. Therefore, restricting $\pi$ induces representations $\overline{\varphi_{E}(e_{i,i})J_{E}\varphi_{E}(e_{i,i})}\to p_{i}\mathbb{B}(\mathcal{H})p_{i}\cong\mathbb{B}(p_{i}\mathcal{H})$, and we will show that these are irreducible. Before proving this, observe that a consequence of this is that, since each $\overline{\varphi_{E}(e_{i,i})J_{E}\varphi_{E}(e_{i,i})}$ is a non-zero and abelian $\mathrm{C}^{\ast}$-algebra (as a subset of $D$), each projection $p_{i}$ must have rank one, and so $\mathcal{H}$ is $r$-dimensional. It follows that $\mathbb{B}(\mathcal{H})$ can be identified with $M_{r}$ in such a way so that $\theta$ becomes the identity map, i.e.
@@
To verify that the restriction of $\pi$ on $\overline{\varphi_{E}(e_{i,i})J_{E}\varphi_{E}(e_{i,i})}$ is irreducible, let $t\in p_{i}\mathbb{B}(\mathcal{H})p_{i}$ be in the commutant of $\pi(\overline{\varphi_{E}(e_{i,i})J_{E}\varphi_{E}(e_{i,i})})$, and set
$\tilde{t}\coloneqq\sum_{j=1}^{r}\theta(e_{j,i})t\theta(e_{i,j})\in\mathbb{B}(\mathcal{H}).$
For any $y\in J_{E}$, we have $\pi(y)=\sum_{j,k=1}^{r}p_{j}\pi(y)p_{k}$ and $\theta(e_{i,j})\pi(y)\theta(e_{k,i})\in\pi(\overline{\varphi_{E}(e_{i,i})J_{E}\varphi_{E}(e_{i,i})})$. A direct calculation using these facts shows that $\tilde{t}$ commutes with $\pi(J_{E})$. Since $\pi$ is irreducible, $\tilde{t}\in\mathbb{C}\mathrm{id}_{\mathcal{H}}$, and thus $t=p_{i}\tilde{t}p_{i}\in\mathbb{C}p_{i}$, as we wanted.
Let $\sigma\colon\overline{\varphi_{E}(e_{i,i})J_{E}\varphi_{E}(e_{i,i})}\to\mathbb{C}$ be the character satisfying
$\pi(x)=\sigma(x)\cdot e_{1,1}$
for all $x\in\overline{\varphi_{E}(e_{i,i})J_{E}\varphi_{E}(e_{i,i})}$. Since $\overline{\varphi_{E}(e_{i,i})J_{E}\varphi_{E}(e_{i,i})}\subset D$, we can extend $\sigma$ to a character on $D$ which we denote again by $\sigma$, and which we can in turn extend to a state on $A$ as $\tilde{\sigma}:=\sigma\circ\Phi$. Define a map $\varrho\colon A\to M_{r}$ as
@@
We claim that, for $g\in D_{+}$ and $e_{i,j}\in M_{r}$ we have $\varrho(g\varphi_{E}(e_{i,j})g)=\pi(g\varphi_{E}(e_{i,j})g)$. Indeed, using that $\varphi_{E}(D_{r})$ is a subset of $D$ which is abelian, by order zero functional calculus we see that $g\varphi_{E}(e_{i,j})g\in\overline{\varphi_{E}(e_{i,i})A\varphi_{E}(e_{j,j})}$ and therefore
@@
which in conjunction with (4.7) shows that $\varrho(g\varphi_{E}(e_{i,j})g)=\pi(g\varphi_{E}(e_{i,j})g)$ as we wanted. It follows now that $\varrho(h_{E})=\pi(h_{E})$. Since $\pi$ was picked so that $\|h_{E}\|=\|\pi(h_{E})\|$, we have that $\|h_{E}\|=\|\varrho(h_{E})\|$. Let $h^{\prime}_{E}:=f^{\prime}a^{1/2}\varphi_{E}^{(j_{0})}\psi_{E}^{(j_{0})}(1_{A})a^{1/2}f^{\prime}$. Since $\|h_{E}-h^{\prime}_{E}\|<\frac{\eta}{2(d+1)}$ by (vi) of Proposition 4.1 and since $\varrho$ is contractive, we have that
@@
Using in the first line of the chain of inequalities below that $\varrho$ is positive and $0\leq h^{\prime}_{E}\leq f^{\prime}af^{\prime}$ (as $\varphi\psi$ is contractive), we also have that